Hi all, I have a bit of a problem with a Matrices question I'm trying to answer.
I need to find the product of the following matrices:
A=
2 -1 2
-1 2 0
0 0 2
2 0 -3

B=
2 -1 2
1 2 0
0 0 2

I also need to state which of the two products should be taken -A . B or B . A.
Can we sum up and A and B?
Any help would be great.
-Nick

Answer by stanbon(48568)

find the product of the following matrices:
A=
2 -1 2
-1 2 0
0 0 2
2 0 -3
---
A is a 4 by 3 matrix.
--------------------------
B=
2 -1 2
1 2 0
0 0 2
---
B is a 3 by 3 matrix
--------------------------
I also need to state which of the two products should be taken -A . B or B . A.
A*B works because the column count on A equals the row count on B.
------------------------
Can we sum up and A and B?
No, because the dimensions are not the same.
-------------------------------------------------------
A=
2 -1 2
-1 2 0
0 0 2
2 0 -3

B=
2 -1 2
1 2 0
0 0 2

AB =
3 -4 4
0 5 -2
0 0 4
4 -2 -2
====================
Cheers,
Stan H.